OpenStudy (anonymous):

Derivative of f(x) = (e^x)^-1

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

sub `u` for `e^x`

OpenStudy (anonymous):

f(x) = u^-1 f'(x) = -1/u^2 I got -1/ e^x . Is that correct?

OpenStudy (solomonzelman):

Yes, but just 1/e^x without the minus

OpenStudy (anonymous):

Please enlighten me?

OpenStudy (solomonzelman):

nvm, you are right... sorry

OpenStudy (solomonzelman):

You used to power rule correctly, I am an idiot -:(

OpenStudy (anonymous):

derivative of e^x = 1(e^x)... oh the weird thing is when I type the answer to the problem it says wrong xD I guess I need parentheses !

OpenStudy (anonymous):

No, you're not! You're smart and awesome. Thank you so much for coming to help me! <3

OpenStudy (solomonzelman):

Anytime... I almost misled you, so be careful with me next time:)

OpenStudy (anonymous):

Aww I'll pay lotsa attention to you xD <3

OpenStudy (solomonzelman):

to me or to my solutions? that makes a difference ..... I know what you mean though :P

OpenStudy (anonymous):

Can't it be both? >.o

OpenStudy (solomonzelman):

Well, perhaps can be so...

OpenStudy (anonymous):

Well it's admiration xP No worries!... yet

OpenStudy (anonymous):

Really nice chatting with you. I'll see you again soon!

OpenStudy (anonymous):

Thanks for helping me out :)

OpenStudy (solomonzelman):

Yes, hopefully we'll meet, yw:)

OpenStudy (anonymous):

definitely will

OpenStudy (solomonzelman):

:)